OpenStudy (albert0898):

Describe Newton’s First Law of Motion. What can a force do to an object? Why do objects eventually stop moving?

OpenStudy (anonymous1):

I'm not an English native speaker, but I'll try to express myself in English. Newton's First Law of Motion states that, if there is no net force acting on an object, the object tends to maintain its state of equillibrium. There are two kinds of equillibrium: static (the object is at rest) and dynamic (the object is moving with a constant speed). In other words, Newton's First Law of Motion states that an object tends to stay at rest or in motion with a constant speed, if there is no net force acting on the object.

OpenStudy (anonymous1):

A force will produce acceleration on the object. This acceleration is proportional to the mass of the object. Suppose an object of mass m. A force of magnitude F will produce on the object an acceleration a=F/m. This is Newton's Second Law: a=F/m, or F=ma. Acceleration means rate of change in speed. When we say that a force produces acceleration on an object, we mean that the force makes the object's speed increase. If the force is constant, then the acceleration is constant; then, the object's speed will change linearly with time. If the object is at rest, and a constant net force starts to act on that object, it will experience a change in it's speed: it will start moving with its speed changing at a constant rate.

OpenStudy (anonymous1):

In practice, objects tend to eventually stop moving, and this generally happens due to friction. Friction is a force that opposes the motion of an object when it moves on a surface. Friction produces a "negative" acceleration on the object, making its speed decrease until it becomes zero.

OpenStudy (anonymous):

the statement " the specific latent heat of fusion of ice is 3.3\times 10^{5}J/kg" means? option 3.3 x 105J of heat energy is require to rise the temperature if ice to melting point . 3.3 x 105 J of heat is absorbed to condense water to ice at 0^{o}c 3.3 x 105 J of heat is required to change 1kg of pure ice at 0^{o}c to 1kg of water at 0^{o}c 3.3 x 105J of heat energy is required to change any amount of pure melting ice at 0^{o}c to water at the same temperature.

OpenStudy (albert0898):

Thanks!